# CushionCatch: Compliant Catching Mechanism for Mobile Manipulators via Combined Optimization and Learning

### v1

@@ VI CONCLUSIONS
We present a framework integrating optimization and learning to enable compliant catching with a cushioning mechanism for mobile manipulators. The framework employs a two-level motion optimization: a high-level capture planner to determine the robot’s optimal catching state, and a low-level joint planner with PRC and POC planner to generate rapid motions and track compliant sequences while ensuring safety, respectively. Compliant control strategies are learned through a P-LSTM network from human demonstrations, which leverage the strengths of LSTM for temporal dependencies and positional encoding for spatial context. Experimental results indicate a 98.70% success rate in simulations and 92.59% in real-world tests, alongside a 33.2% decrease in impact torque. Moreover, due to the modular design of the CCCM framework, we envision that it can be easily ported to other robots or fields.

### v3

@@ VI CONCLUSIONS
This paper aims to address the challenge of compliant catching for mobile manipulators. We have developed a novel CCMM framework that combines optimization and learning to address all the planning and tracking challenges during the catching process. A high-level planner is used to calculate the optimal capture point and joint configuration, while a PRC planner ensures the manipulator reaches the target joint configuration as quickly as possible. The compliant strategies are learned from human demonstrations, with a P-LSTM network proposed to capture temporal dependencies and spatial context through positional encoding. Additionally, a POC planner is proposed to track the sequence predicted by the network while addressing safety concerns arising from structural differences between humans and robots. Experimental results show a 98.70% success rate in simulations, 92.59% in real-world tests, and a 28.7% reduction in impact torque. The limitation of our framework is its reliance on the accuracy of external sensing, as the data from the initial frames is crucial for subsequent predictions and planning. Moving forward, a key focus will be exploring how to complete above tasks using only the robot’s camera. Finally, due to the modular design of our CCMM framework, we envision that it would be easily ported to other robots or fields.